# Security Approaches for Data Provenance in the Internet of Things: A Systematic Literature Review

### v1

@@ 5. Security techniques for data provenance
5.8   Path Difference — Path difference is the sum of an indicator variable which represent whether the actual packet path for the next hop is the same as the parent node of this packet along the routing path. Parent information can be used in reference packets for path reconstruction if the path difference value is equal to zero. If not, more time will be required to record the real forwarders. In order to minimize the message overhead, it is better to have a minimal path difference value. In big-scale sensor networks having lossy links and complex routing dynamics, Gao et al. (2013) mention the ineffectiveness of current path reconstruction techniques. The authors present Pathfinder, a cutting-edge path reconstruction method. Pathfinder takes advantage of temporal correlation between a group of packet paths at the node side and effectively compresses the route information with path variation; at the PC side, Pathfinder determines packet routes from the compressed data and uses smart path speculation to reassemble the data packet paths with a high reconstruction ratio. With the use of comprehensive simulations and traces from a large real-world sensor network, they construct Pathfinder and evaluate its performance against the two most similar approaches. Results indicate that Pathfinder performs noticeably better than MNT (Keller et al., 2012) and PathZip (Lu et al., 2012) in a number of network setups. Their findings show that Pathfinder achieve both low transmission overhead and high reconstruction ratio.
@@
Sadineni et al. (2023) propose a template-based provenance approach, ProvLink-IoT, to provide reliable forensic analysis in IoT networks. ProvLink-IoT is developed to analyze link-layer attacks. Many open source tools are used to implement the system in a simulated environment to approve its robustness in correlating evidence that is found in link-layer provenance. Based on provenance logs and network data gathered from the network, traceability graphs are generated in both normal and attack scenarios. The approach is studied using 6TiSCH network stack. To analyze the effects and conduct forensic analysis, the authors applied three link-layer attacks to the Time Slotted Channel Hopping (TSCH) and 6top (i.e. operational sublayer) layers of the 6TiSCH network. For performance evaluation, they used storage overhead and provenance growth rate.
@@
5.10   Frameworks using Storing Methods — Researchers have used various storage methods, using database placement strategies, to develop frameworks designed for diverse IoT applications. These methods include a number of techniques for storing and querying provenance information within the database. They offer management solutions to ensure the secure storage of provenance data while meeting security requirements. These frameworks offer a complete cycle for tracking the history of provenance records, covering everything from initial capture to storage and analysis. In this section, we present selected works that have designed such frameworks for managing data provenance.
@@
After analyzing the selected state-of-the-art papers based on our proposed taxonomy in Figure 3, the findings of this analysis is shown in Tables 4 to  8 in Appendix A. In Table 4, each paper is analyzed based on the provenance encoding method, provenance storage method, application, security analysis, advantages and shortcomings. We present the main performance metrics that are needed to evaluate any data provenance approach in IoT networks that are provenance length, energy comparison, data packet size, link-loss rate, detection rate, false positive rate, false negative rate and computation time. We take into consideration for the analysis of each selected approach these performance metrics which are presented in Table 6. Moreover, in Table 5, we provide details on the provenance category, implementation, and simulation software. This resource is presented to help researchers in exploring the different implementation tools and simulation software associated with each provenance category. Table 7 summarizes the security requirements for data provenance integration in IoT networks, previously introduced in Section 3. Similarly, Table 8 summarizes the security attacks illustrated in Figure 8.
6. Discussion
7. Open issues and research challenges
In order to open up additional discussion and suggest recommendations for enhancing the development of data provenance in the IoT networks, we have outlined several open problems, research gaps, challenges and limitations from the studied literature.
  • Watermarking as a solution: Maintaining data integrity through procedural requirements is the major goal of watermarking system design. These requirements cover node ownership, data integrity, and bandwidth establishment. Embedding a watermark with data is the main technique to achieve the requirements in watermarking schemes (Sultana et al., 2013, 2015; Dai et al., 2008; Hussain et al., 2014; Panah et al., 2015). The criteria used in the generation of watermarking schemes is not restricted to WSN; it also includes the security of multimedia programs (Sultana et al., 2011; Hasan et al., 2009b; Prasad and Koliwad, 2010; Fallahpour and Megías, 2015) and database formation (Jain and Jain, 2015; Gosavi and Mali, 2015; Khan et al., 2013). This wide range of domains allows watermarking to be one of the lightweight solutions in limited networks such as IoT. Hence, the features of data hiding that watermarking provides a new provenance encoding technique to store provenance records in an efficient way in-terms of storage and transmission overhead. Only one relevant work (Sultana et al., 2011) considered watermarking for data provenance. They consider inter-packet timing characteristics for embedding provenance information and does not use the data features and watermarking embedding with the data packets. This technique can be used to develop lightweight solutions for the limitations of IoT networks in terms of storage and computation.
  • Lack of complete provenance management system and narrow focus: The literature is still missing a robust provenance management system that takes into account all required procedures from collecting, storing and analyzing provenance information. The system should satisfy the security requirements while maintaining storage and data management issues. To achieve this, these systems should be designed based on the architecture of IoT networks, taking into account various factors, such as communication protocols, computational capabilities, storage limitations, node coordination, database integration, and potential security attacks against data and provenance records. Additionally, it is essential to thoroughly test the robustness of the proposed provenance system, assessing its performance across all relevant metrics. This approach ensures that the system complies with all the necessary objectives, rather than evaluating its performance based on only a subset of criteria. Many of the existing approaches fail to adequately address these important issues while maintaining the security of provenance data. Furthermore, the proposed provenance approaches focus on a specific application for a specific domain. The fast development of interconnected applications of different domains require adapting these techniques to be applied to multiple domains that include different structure of provenance records.
  • Integrating data provenance with Intrusion Detection System (IDS): Adding data provenance to system logs transforms them from simple records into rich sources of security insights which are used in traditional Host-based Intrusion Detection Systems (HIDS). This helps investigators to detect security threats with greater accuracy, minimizing computational time and resources on false alarms, as extensive research has documented (Ayoade et al., 2020; Barre et al., 2019; Berrada and Cheney, 2019; Xie et al., 2016, 2020; Han et al., 2017; Du et al., 2017; Gao et al., 2018b; Xie et al., 2021; Berrada et al., 2020; Gao et al., 2018a; Garchery and Granitzer, 2020; Hassan et al., 2020; Wang et al., 2020). Data provenance can be integrated with HIDS to improve intrusion detection. IDS based on provenance, known as Provenance-based Intrusion Detection Systems (PIDS), leverage data provenance to identify intrusions. This involves examining not only the properties of system entities but also dealing with the causal relationships and information flow within a provenance graph (Zipperle et al., 2022). However, the integration of data provenance and IDS in IoT networks has not been widely studied. This opens a crucial field for researchers to study, addressing the significance of integrating data provenance in IoT networks and exploring its various applications for enhancing intrusion detection.
8. Conclusion
Appendix A Additional Figures and Comparison Tables
Figure 6. Provenance domains and the scope of our research.
Figure 7. IoT network model assumed in our work.
Figure 8 provides a taxonomy of attacks against IoT networks.
Figure 8. Taxonomy of security attacks in IoT networks with data provenance approach.
Table 4. Comparative analysis of selected studies.
@@ Appendix A Additional Figures and Comparison Tables
Bloom Filters
@@
Fingerprints
@@
Watermarking
Physical Unclonable Functions
[…]
Graph-based
 Chang et al. (2022) {NoHyper} (2022) Event-flow graphs Provenance server IoT smart apps Observes provenance without utilizing a sophisticated query language by just selecting the appropriate nodes on an event flow graph. As the number of events and actions increases, the provenance data becomes large in scale over time, provenance data is not securely stored in the provenance server.
Graph-based
Data Sanization
Lexial Chaining
Path Difference
Logging-based
Frameworks using storing methods

### v2

@@ 7. Security techniques for data provenance
5.4   Path Difference — Path difference is the sum of an indicator variable which represents whether the actual packet path for the next hop is the same as the parent node of this packet along the routing path. Parent information can be used in reference packets for path reconstruction if the path difference value is equal to zero. If not, more time will be required to record the real forwarders. In order to minimize the message overhead, it is better to have a minimal path difference value. In big-scale sensor networks having lossy links and complex routing dynamics, Gao et al. (2013) mention the ineffectiveness of current path reconstruction techniques. The authors present Pathfinder, a cutting-edge path reconstruction method. Pathfinder takes advantage of temporal correlation between a group of packet paths at the node side and effectively compresses the route information with path variation; at the PC side, Pathfinder determines packet routes from the compressed data and uses smart path speculation to reassemble the data packet paths with a high reconstruction ratio. With the use of extensive simulations and traces from a large real-world sensor network, they construct Pathfinder and evaluate its performance against the two most similar approaches. Results indicate that Pathfinder performs noticeably better than MNT (Keller et al., 2012) and PathZip (Lu et al., 2012) in a number of network setups. Their findings show that Pathfinder achieve both low transmission overhead and high reconstruction ratio. An overview of the proposed technique is shown in Table 5.
@@
Sadineni et al. (2023) propose a template-based provenance approach, ProvLink-IoT, to provide reliable forensic analysis in IoT networks, as presented in Table 5. ProvLink-IoT is developed to analyze link-layer attacks. Many open source tools are used to implement the system in a simulated environment to approve its robustness in correlating evidence that is found in link-layer provenance. Based on provenance logs and network data gathered from the network, traceability graphs are generated in both normal and attack scenarios. The approach is studied using 6TiSCH network stack. To analyze the effects and conduct forensic analysis, the authors applied three link-layer attacks to the Time Slotted Channel Hopping (TSCH) and 6top (i.e. operational sublayer) layers of the 6TiSCH network. For performance evaluation, they used storage overhead and provenance growth rate.
Table 5. Overview of selected studies using Watermarking, Sanitization, Lexical Chaining, Path Difference and Logging-based methods for provenance encoding in IoT networks.
Method Method Analysis
Watermarking
Data Sanitization
Lexial Chaining
Path Difference
Logging-based
5.6   Bloom Filters — A compact data structure, using hashing, enables rapid verification of item presence within the structure. Provenance information is embedded within the generated structure and the bloom filter is transmitted along with the data. By employing this method, the original information remains inaccessible to potential adversaries. Due to the variability of the bloom filter value from one packet to another, establishing a connection or association between previous bloom filters and recent data becomes challenging. An overview of the selected papers in this category are shown in Table 6. In their work,  Sultana et al. (2015) present a secure provenance scheme for WSN. Their approach involves embedding provenance information into a Bloom Filter, which is transmitted alongside the data. This scheme effectively addresses the challenges posed by resource constraints in WSN. It requires a single channel for data and provenance transmission. The scheme overlooks data integrity and only focuses on studying the packet drop attack. Additionally,  Siddiqui et al. (2019) present a data provenance technique for IoT devices that employs Bloom Filter and attribute-based encryption. This approach presents challenges as IoT devices typically have limited memory capacity, making it impractical to store extensive provenance information. Furthermore, the technique is vulnerable to physical attacks, whereby an attacker can easily manipulate the stored provenance information within an IoT device.
Harshan et al. (2020) introduce a method for embedding provenance information known as the Deterministic Double-Edge (DDE) embedding mechanism, which is based on Bloom filters (Lu et al., 2012; Shebaro et al., 2012). In this approach, a relay node embeds information about both edges connected to it: the edge through which it receives a packet and the one through which it sends the packet. This means that a node can effectively cover two edges in the path with a single action, reducing the need for the next node in the path to update the provenance. This results in decreased packet delay. Authors show that the hop-counter within the provenance segment can also be used to coordinate the skipping strategy among nodes. Additionally, they propose upper bounds on the error rates of the DDE embedding technique, which are expressed as functions of the network’s node count, the number of hops, the size of the Bloom filter, and the number of hash functions computed by each node.
Table 6. Overview of selected studies using bloom filters-based methods for provenance encoding in IoT networks.
Method Method Analysis
Bloom Filters
[…]
5.7   Fingerprints — Fingerprints refers to unique identifiers or signatures assigned to the communication links between IoT devices. These fingerprints are typically generated based on various characteristics or attributes of the link, such as received signal strength, latency, packet loss, or other network metrics. The homomorphic feature of public-key cryptography is exploited by numerous anonymous fingerprinting systems. These approaches enable the user’s fingerprint to be embedded in an encrypted domain (using public key) so that only the user may access the decrypted fingerprinted material by using the private key (Megías, 2015). Data provenance solutions have been developed by analyzing and comparing link fingerprints generated from different attributes. The selected studies based on this technique is shown in Table 7. In their work,  Ali et al. (2014) present a method for enhancing the security of data provenance in bodyworn medical sensor devices. They achieve this by using the spatio-temporal characteristics of the wireless channels used for communication by these devices. This solution allows two parties to create highly similar link fingerprints, which uniquely link a data session to a specific wireless connection. This link enables a third party to later verify transaction details, especially the specific wireless link through which the data was transmitted. To validate this approach, experimental testing was conducted using MicaZ motes running TinyOS, and the results show an improvement in energy efficiency. Additionally, the proposed technique generates provenance information on each session, reducing the use of cryptographic methods. Using different fingerprint method, Alam and Fahmy (Alam and Fahmy, 2014) propose a provenance encoding and construction method that adapts three encoding schemes: juxtaposition of ranks, prime scheme, and Rabin fingerprints. The method is referred to as Probabilistic Packet Flow (PPF). In the first scheme, provenance is constructed based on the rank of the node instead of the node ID, which requires fewer bits for encoding. Assuming that the packet contains a specific space to hold the identities of $m$ nodes, a counter of $\log(m)$ bits is used to track the embedded ranks in the packet. The second scheme is based on prime multiplication, which embeds more node IDs using the same number of bits compared to the previous one. This method uses prime numbers as node IDs, and their multiplication results in encoding a set of IDs that can be uniquely factorized. In their third method, a data path traversed by a packet is a sequence of bits that represents the IDs of nodes on that path. The fingerprint of this sequence of bits is transmitted instead of the actual sequence.  Kamal and Tariq (2018) introduce a lightweight protocol for a multi-hop IoT network, aiming to ensure both data security and the establishment of data provenance. The protocol uses link fingerprints derived from the Received Signal Strength Indicator (RSSI) of IoT nodes within the network. The protocol achieves data provenance by appending the encoded link fingerprint to the data packet header as it traverses each node. After receiving the packet, the server decode the packet header in a sequential process. However, provenance data expands rapidly, which requires transmitting a large amount of provenance information to data packets, thereby increasing the bandwidth overheads.
Table 7. Overview of selected studies using fingerprints-based methods for provenance encoding in IoT networks.
Method Method Analysis
Fingerprints
@@
5.8   Frameworks using Storing Methods — Researchers have used various storage methods, using database placement strategies, to develop frameworks designed for diverse IoT applications. These methods include a number of techniques for storing and querying provenance information within the database. They offer management solutions to ensure the secure storage of provenance data while meeting security requirements. These frameworks offer a complete cycle for tracking the history of provenance records, covering everything from initial capture to storage and analysis. In this section, we present selected works that have designed such frameworks for managing data provenance and provide an overview of these works in Table 8.
@@
Table 8. Overview of selected studies based on frameworks with storing methods for provenance encoding in IoT networks.
Method Method Analysis
Frameworks using storing methods
5.9   Graph-based Provenance — Graph-based provenance is a method used in data provenance encoding to represent the origin or history of data. These graphs capture the relationships and dependencies among data elements, processes, and transformations, providing a visual or structural representation of how data has evolved over time. Three different types of graph-based provenance representations as descibed below and an overview of these methods is provided in Table 9.
Event-flow graphs are a type of provenance graph that represents data provenance by capturing events and the flow of data between them. Each event corresponds to a data operation or transformation, and edges in the graph indicate the data flow from one event to another.  Chang et al. (2022) propose a data provenance approach for provenance systems in IoT applications based on event-action flows instead of provenance graphs. Event-action flow is defined as a sequence of events and actions including a time-stamp in an execution trace. They are less complex than provenance graphs due their simple structure form which allows users to understand it easier. In their work, the authors present an event-flow graph to regular users as a static abstraction of every possible provenance graph for IoT applications. They dynamically link time-stamped events and actions to statically create event and action nodes. Then, users can query provenance information from the event-flow graph by selecting an event or action node to choose the associated time-stamped actions or events. After users create a query by picking a timestamped event or action, the system will respond with one of two types of information: “What provenance” answers questions about which events or actions are triggered by the user’s specified action or event, and “Why provenance” provides insights into the events or actions responsible for causing the user-specified action or event. The system is developed as a Graphical User Interface (GUI) forming a user-friendly graphical provenance system.
Finally, the use of Directed Acyclic Graphs (DAG) allows directed structures with no cycles. This means that data transformations and dependencies are represented as a directed flow without loops or feedback. DAGs are used to represent provenance information because they are well-structured and provide efficient querying and tracking of provenance records. ProChain framework is a provenance-aware approach of traceability proposed by Al-Rakhami and Al-Mashari (Al-Rakhami and Al-Mashari, 2022) for IoT-based supply chain systems. The IOTA protocol, a third-generation DLT, is used by the ProChain framework. To overcome limitations and provide a scalable, quantum-resistant, and attack-proof solution for the systems built around IoT, it makes use of the DAG information structure in contrast to the linear structure used by the blockchain (Silvano and Marcelino, 2020). ProChain enables food item traceability from manufacturing to retailer with the use of several IoT sensors and provenance data at each engaged supply chain phase. ProChain strengthens and improves the management and optimization of all operations while also serving as a guarantee for the quality and safety of food. On the Raspberry Pi 3B platform, the ProChain idea is evaluated by simulating an IoT-deployed supply chain. The average measured time and energy consumption were then evaluated to check on the usability of the framework. The authors review and implement the framework to show how it may be used in supply chain systems to couple supply chain data with the IOTA Tangle and generate provenance information by adding data for various payload sizes.
Table 9. Overview of selected studies using graph-based techniques for provenance encoding in IoT networks.
Method Method Analysis
Graph-based
 Chang et al. (2022) {NoHyper} (2022) Event-flow graphs Provenance server IoT smart apps Observes provenance without using a sophisticated query language by just selecting the appropriate nodes on an event flow graph. As the number of events and actions increases, the provenance data becomes large in scale over time, provenance data is not securely stored in the provenance server.
5.10   Cryptography-based Techniques — Various cryptographic techniques, including symmetric and asymmetric cryptography methods, hash functions, and digital signatures, are applied to design and implement systems for identifying the source of data and ensuring the integrity of both data and provenance. These concepts are used to propose solutions for data provenance in many applications. The literature includes the largest number of these solutions when dealing with cryptographic approaches as shown in Table 10. The trustworthiness of the data items in a sensor network is evaluated by  Lim et al. (2010). They compute trust scores for sensor nodes and data packets by using data provenance. These scores provide a method of indicating the level of trustworthiness of both nodes and data items. This method provide security and trustworthiness for sensor networks, yet they fail to address the challenge of retrieving data provenance against various attacks. Moreover, a dictionary-based secure provenance approach for WSN is proposed by  Wang et al. (2016a). They embed path indexes in the provenance instead of the actual data path by using packet path dictionaries. Therefore, compared to the current lossy provenance systems, the compressed provenance size in the proposed lossless approach is smaller. To achieve security requirements such as integrity, availability and authenticity of provenance, the AM-FM sketch method and a robust packet sequence number generating technique are used in the system. Provenance records are encoded at nodes that engage in every stage of data processing and transmission using the suggested dictionary-based system. A secure message authentication code integrates the packet and its provenance together to provide security against any unauthorized change. After verifying the Message Authentication Code (MAC) during decoding, the BS retrieves the packet’s provenance graph.
While the majority of previous studies have concentrated on how to protect against data manipulation in Home Area Networks (HAN) networks,  Chia et al. (2017) introduce a security topic that has received less attention in such networks, namely data provenance. To ensure that the stated energy usage is actually consumed and is gathered from the specified node at the exact location, they provide a unique technique based on threshold cryptography and Shamir’s secret sharing (Shamir, 1979). The authors describe a unique use of secret sharing schemes for home energy monitoring networks, in which a secret, or a single private key, is distributed among all the members who provide data on energy consumption. This system also addresses the issue of location verification. To achieve this, authors incorporate a location generator as an additional component. The location generator employs trilateration techniques based on RSSI values and RSSI filtering methods to verify the location of the device. This not only ensures that the right power source is under observation but also detects any potential device relocation. Also,  Suhail et al. (2018) provide a solution to the challenge of integrating IoT with a system that is aware of data provenance, enabling the tracking of data flow across nodes and monitoring data transformations applied by nodes in the network. They introduce a lightweight method for transmitting provenance information for IoT sensor data. This method encodes data provenance information using a hash chain scheme as it traverses each participating node, with the final provenance verification taking place at the destination node. The technique is implemented within the context of the IPv6 Routing Protocol for Low-Power and Lossy Networks (RPL) (Alexander et al., 2012) in Contiki OS (Dunkels et al., 2004). Each generated data packet includes three fields: packet sequence number, data payload and provenance information (i.e. hash value of node ID). Furthermore,  Xu et al. (2019) present a provenance approach for WSN based on path index differences, where a new packet path is encoded using the index of a highly similar backbone path and the differences between them. Even if the topology of the WSN is unstable, this technique may reach an increased provenance compression rate compared to the dictionary based provenance scheme (Wang et al., 2016a). In the presented backbone path selection approach for WSN, the gradient of the node on the data packet path is also determined. In order to cover the commonly used packet transmission channels, the authors also develop a comparable backbone path elimination technique for WSN. To find the packet pathways with the highest similarity in the dictionaries, they develop a locality-sensitive hashing (LSH), SimHash (Buyrukbilen and Bakiras, 2014; Charikar, 2002), based similarity comparison algorithm. The proposed approach is evaluated according to Total Energy Consumption (TEC) and Average Provenance Size (APS).
In their research,  Xu et al. (2022) introduce two provenance schemes for Wireless Sensor Networks. The first is the Path Index Differences-based Provenance (PIDP) scheme, where provenance information is encoded using differences in packet path indices. Specifically, it links an index representing the packet path’s similarity to the main backbone path and the variation between the actual packet path and this backbone path. Additionally, the authors propose a second scheme known as the Path Hash Value-based Provenance (PHP) scheme. In this approach, provenance data is encoded as a combination of the data source node’s ID and a part of the hash value taken from the packet’s path. Both schemes are evaluated based on their provenance compression rate and energy conservation rate, and these metrics are then compared to those of the DP scheme. Xu and Wang (Xu and Wang, 2022) introduce a provenance scheme called Multi granularity Graphs-based Stepwise Refinement Provenance (MSRP). In this scheme, they use mutual information between pairs of nodes as a similarity index to categorize node IDs. This categorization forms the basis for generating multi granularity topology graphs. Additionally, they apply the Dictionary-based Provenance (DP) scheme for stepwise encoding of the provenance information. The Base Station (BS) follows the same stepwise approach for provenance recovery and simultaneously conducts data trustworthiness evaluation during decoding. The performance of the MSRP scheme is thoroughly evaluated through a combination of simulations and testbed experiments. Results show that the scheme achieve high provenance compression rate, energy usage and efficiency of data trustworthiness assessment.
Table 10. Overview of selected studies using cryptography-based methods for provenance encoding in IoT networks.
@@
5.11   Physical Unclonable Functions — Physical Unclonable Functions (physical unclonable functions (PUFs)) provide a hardware-oriented system that generates a response for a specific challenge, ensuring uniqueness for each device in the system. PUFs facilitate the authentication of IoT devices that generate sensitive data while maintaining device anonymity. Through the presence of a trusted third-party verifier, the identities of these devices can be verified without compromising their anonymity (Kanuparthi et al., 2013). The selected techniques are shown in Table 11. Aman et al. (2019) developed two secure protocols for data provenance in IoT networks, aiming to achieve authentication and privacy preservation. These protocols address two different scenarios: the first scenario involves a direct connection between an IoT device and a wireless gateway, while the second scenario deals with IoT devices indirectly connected to the wireless gateway through multiple hops of other IoT devices. Both protocols use PUFs in addition to wireless link fingerprints generated from the RSSI between communicating nodes. Through experimental analysis, the authors show that these protocols achieve high efficiency in terms of computational complexity and energy usage, while also obtaining robustness against various physical and cloning attacks. Aman et al. (2021) propose an analytical model to create a mechanism that enables the establishment of data provenance in IoT systems. Their approach incorporates PUFs and the extraction of fingerprints from the wireless channel, along with the implementation of mutual authentication and anonymity measures, all aimed at achieving robust data provenance. The approach lacks consideration for the multi-hop scenario and fails to adequately address tracking of data packet provenance. Moreover, Hamadeh and Tyagi (Hamadeh and Tyagi, 2021) propose an approach which combine two solutions: data provenance and privacy prserving in IoT networks. They provide trustworthiness and dependable IoT networks by using PUFs with non-interactive zero knowledge proof. In this method, an IoT device has the capability to transmit data to its respective server without revealing its identity, as it provides proof of ownership. In particular, the method under consideration is designed to validate that the authorized device executed an authorized program for creating or modifying data. Authors introduce a privacy-centric data provenance protocol. To validate practicality and effieciency, they developed the protocol using Altera Quartus and subsequently implemented it on an Altera Cyclone IV FPGA.
Table 11. Overview of selected studies using watermarking and Physical Unclonable Functions for provenance encoding in IoT networks.
Method Method Analysis
Physical Unclonable Functions
@@
[…]
5.12   Blockchain-based Solutions — Distributed ledger technologies, like blockchain, have gained visibility for maintaining the security and privacy of provenance data. These blockchain platforms serve as the primary means to ensure data integrity by creating transactions as provenance records that are chained in a secure manner. Once recorded on the blockchain, these records cannot be altered or removed, thereby establishing a trustworthy system for verifying the integrity of the provenance information in a network environment that is vulnerable to different attacks. Table 12 provides an overview of the selected studies that uses blockchain-based methods to maintain the security of provenance data in IoT networks.
In their work,  Javaid et al. (2018) propose a solution for data integrity and data provenance in IoT networks by using PUFs and a blockchain variant with smart contracts that is Ethereum. This method is called BlockPro. PUFs are used to maintain data provenance by providing unique hardware fingerprints. To overcome data tampering attacks and block unregistered devices, Ethereum is used as a decentralized digital ledger. With the presence of expanding sequence of records, data undergoes initial validation before being permanently stored on the blockchain. Once stored, it can not be tampered with or modified, thus ensuring data integrity.  Sigwart et al. (2020) implement an IoT data provenance framework based on smart contracts using a generic data model to provide data provenance capturing, storing, and querying functionalities for different IoT use cases. This work extends their approach proposed in (Sigwart et al., 2019). The framework uses the data provenance model by  Olufowobi et al. (2017). The authors conduct an assessment of the proposed framework with specified requirements through the implementation of a proof-of-concept using Ethereum smart contracts. In this approach, a provenance record ($prov(dp)$) for a data point ($dp$) is composed of a 3-part structure. It links the address ($addr(dp)$) of $dp$ (i.e. essentially an identifier ID) with the set of provenance records associated with the data points that are used to create a $dp$ (referred to as $inputs(dp)$), and includes a context element ($context(dp)$). This element includes information for provenance purposes, such as details about the agents involved in the computation of the data point, timestamp, location, or the specific execution context within the IoT system.
In their work, Porkodi and Kesavaraja (Porkodi and Kesavaraja, 2021) present a framework to secure the data provenance by introducing blockchain and access control policies in IoT systems. They use hybrid attribute-based encryption to securely transmit provenance information. The proposed method is evaluated using different performance metrics such as computational cost and the throughput of encryption/decryption, and the key strength is calculated using the effect of avalanche. Additionally, the authors conducted experiments to prove that the proposed approach reduces computational cost and achieves high throughput.
Table 12. Overview of selected studies using blockchain-based methods for provenance encoding in IoT networks.
Method Method Analysis
@@
After analyzing the selected state-of-the-art papers based on our proposed taxonomy in Figure 3, the findings of this analysis are shown in Tables 10 to  8 and Tables 13 to  16 in Appendix A. Each paper is analyzed based on the provenance encoding method, provenance storage method, application, security analysis, advantages and shortcomings. The findings from this in-depth analysis of the presented literature is shown in details in the Discussion section below.
8. Discussion
Based on the insights gained from the analysis and in-depth study of the security techniques for data provenance presented in Section 7, we can better understand the critical challenges and considerations outlined above. These security techniques not only address the issues of trust and integrity within data provenance but also inform our recommendations for enhancing the development and implementation of effective provenance solutions in IoT networks.
In Table 13, we provide details on the evaluation mechanisms used by the various techniques, categorizing them as hardware-based, cloud-based, or simulation-based (software-based or simulator). This resource aims to assist researchers in exploring the different implementation tools and simulation software associated with each provenance category. Our analysis reveals that the majority of studies predominantly use simulation-based approaches to implement and test their techniques. These software-based simulators offer flexibility and cost-effectiveness, making them a popular choice among researchers. However, only a limited number of studies combine both hardware and simulator implementations, which could provide a more thorough evaluation of the techniques in real-world scenarios. This observation highlights a potential gap in the literature, where practical, hardware-based testing could validate the effectiveness of proposed solutions under more realistic conditions. Encouraging a balance between simulation and hardware implementations may lead to more robust and applicable techniques for data provenance in IoT environments.
Moreover, in Table 14, we present the performance metrics evaluated by various security techniques for data provenance in IoT, based on a selection of recent literature. Our analysis reveals that while several metrics are available for assessing the effectiveness of provenance mechanisms—such as provenance length, energy consumption, data packet size, link-loss rate, detection rate, false positive rate, false negative rate, and computation time—most existing studies focus primarily on a limited subset of these, typically provenance length, energy consumption, and computation time. This selective focus indicates that while these metrics are essential, there is limited research that thoroughly evaluates techniques across multiple performance dimensions. For instance, metrics like detection rate, false positive rate, and false negative rate are important for assessing the security and reliability of provenance techniques. However, few studies include these alongside resource-related metrics, such as energy consumption and computation time. Similarly, link-loss rate, which can impact communication stability in IoT networks, is often overlooked. By examining Table 14, we observe a need for more holistic evaluations that consider the full spectrum of performance metrics. Expanding future work to address this broader set of metrics could provide more robust assessments of the practical feasibility of these techniques. This would also align with IoT requirements, where both resource efficiency and high security are critical. The lack of studies that comprehensively address multiple performance metrics highlights an area for further research in developing balanced, resource-conscious, and security-oriented solutions for IoT provenance.
9. Research challenges and open issues
Following our analysis of the security techniques presented in Section 7 and the discussions in Section 8, we aim to foster further dialogue and propose recommendations for advancing data provenance in IoT networks by outlining several open problems, research gaps, challenges, and limitations identified in the reviewed literature.
  • Watermarking as a solution: Maintaining data integrity through procedural requirements is the major goal of watermarking system design. These requirements cover node ownership, data integrity, and bandwidth establishment. Embedding a watermark with data is the main technique to achieve the requirements in watermarking schemes (Sultana et al., 2013, 2015; Dai et al., 2008; Hussain et al., 2014; Panah et al., 2015). The criteria used in the generation of watermarking schemes is not restricted to WSN; it also includes the security of multimedia programs (Sultana et al., 2011; Hasan et al., 2009b; Prasad and Koliwad, 2010; Fallahpour and Megías, 2015) and database formation (Jain and Jain, 2015; Chhaya S. Gosavi, 2015; Khan et al., 2013). This wide range of domains allows watermarking to be one of the lightweight solutions in limited networks such as IoT. Hence, the features of data hiding that watermarking provides a new provenance encoding technique to store provenance records in an efficient way in-terms of storage and transmission overhead. Only one relevant work (Sultana et al., 2011) considered watermarking for data provenance. They consider inter-packet timing characteristics for embedding provenance information and does not use the data features and watermarking embedding with the data packets. This technique can be used to develop lightweight solutions for the limitations of IoT networks in terms of storage and computation.
  • Lack of complete provenance management system and narrow focus: The literature is still missing a robust provenance management system that take into account all required procedures from collecting, storing and analyzing provenance information. The system should satisfy the security requirements while maintaining storage and data management issues. To achieve this, these systems should be designed based on the architecture of IoT networks, taking into account various factors, such as communication protocols, computational capabilities, storage limitations, node coordination, database integration, and potential security attacks against data and provenance records. Additionally, it is essential to thoroughly test the robustness of the proposed provenance system, assessing its performance across all relevant metrics. This approach ensures that the system complies with all the necessary objectives, rather than evaluating its performance based on only a subset of criteria. Many of the existing approaches fail to adequately address these important issues while maintaining the security of provenance data. Furthermore, the proposed provenance approaches focus on a specific application for a specific domain. The fast development of interconnected applications of different domains require adapting these techniques to be applied to multiple domains that include different structure of provenance records.
  • Integrating data provenance with Intrusion Detection System (IDS): Adding data provenance to system logs transforms them from simple records into rich sources of security insights which are used in traditional Host-based Intrusion Detection Systems (HIDS). This helps investigators to detect security threats with greater accuracy, minimizing computational time and resources on false alarms, as extensive research has documented (Ayoade et al., 2020; Barre et al., 2019; Berrada and Cheney, 2019; Xie et al., 2016, 2020; Han et al., 2017; Du et al., 2017; Gao et al., 2018b; Xie et al., 2021; Berrada et al., 2020; Gao et al., 2018a; Garchery and Granitzer, 2020; Hassan et al., 2020; Wang et al., 2020). Data provenance can be integrated with HIDS to improve intrusion detection. IDS based on provenance, known as Provenance-based Intrusion Detection Systems (PIDS), leverage data provenance to identify intrusions. This involves examining not only the properties of system entities but also dealing with the causal relationships and information flow within a provenance graph (Zipperle et al., 2022). However, the integration of data provenance and IDS in IoT networks has not been widely studied. This opens an important field for researchers to study, addressing the significance of integrating data provenance in IoT networks and exploring its various applications for enhancing intrusion detection.
10. Conclusion
Appendix A Additional Figures and Comparison Tables
Figure 6. Provenance domains and the scope of our research.
Figure 7. IoT network model assumed in our work.
Figure 8 provides a taxonomy of the usage and applications for security techniques in IoT networks.
Figure 8. Taxonomy of the usage and applications for security techniques in IoT networks.